A condidate was to submit 15 from a certain number but mistakenly added 25 and her answer was 145. What is the percentage error
Found 2 solutions by ewatrrr, ikleyn:
Answer by ewatrrr(24785)


Hi
A candidate was to 'submit 15 from' a certain number 'x' but mistakenly added 25
and her answer was 145.
 x + 25 = 145 
 x = 120
120 - 15  = 105 ***   | 'submit 15 from'
  105 the correct amount
What is the percentage error:
( (145-105)/105)*100 = 38% is the percentage error (to nearest whole number
Wish You the Best in your Studies.


Answer by ikleyn(52847)

.

                Couple of notices regarding using  English words.


Instead of  " condidate "  write  " candidate ".

Instead of  " to submit 15 from a certain number "  write  "to subtract 15 from a certain number".


Learn  English parallel to  Math from this forum.
Have doubled benefits  ( ! )